A (1,6) , B (-1,4) , C (2,1) are the vertices of a traingle.Show by using distance formula that the triangle ABC is rightangled.
Answer by mananth(16946)

A (1,6) , B (-1,4) , C (2,1) are the vertices of a traingle.Show by using distance formula that the triangle ABC is rightangled.
AB ^2 = (6-4)^2+(1-(-1))^2
=2^2+2^2
8
BC = (4-1)^2+(-1-2)^2
=3^2+(-3)^2
=9+9
=18
CA^2=(6-1)^2+(2-1)^2
=5^2+1^2
=25+1
=26
8+18=26
AB^2+BC^2= AC^2
Hence it is a right triangle